### v1

# Where Should I Study? Biased Language Models Decide! Evaluating Fairness in LMs for Academic Recommendations

Abstract
Large Language Models (LLMs) are increasingly used as daily recommendation systems for tasks like education planning, yet their recommendations risk perpetuating societal biases. This paper empirically examines geographic, demographic, and economic biases in university and program suggestions from three open-source LLMs: LLaMA-3.1-8B, Gemma-7B, and Mistral-7B. Using 360 simulated user profiles varying by gender, nationality, and economic status, we analyze over 25,000 recommendations. Results show strong biases: institutions in the Global North are disproportionately favored, recommendations often reinforce gender stereotypes, and institutional repetition is prevalent. While LLaMA-3.1 achieves the highest diversity, recommending 481 unique universities across 58 countries, systemic disparities persist. To quantify these issues, we propose a novel, multi-dimensional evaluation framework that goes beyond accuracy by measuring demographic and geographic representation. Our findings highlight the urgent need for bias consideration in educational LMs to ensure equitable global access to higher education.
Where Should I Study? Biased Language Models Decide!
Evaluating Fairness in LMs for Academic Recommendations
Krithi Shailya*, Akhilesh Kumar Mishra, Gokul S Krishnan+, Balaraman Ravindran Centre for Responsible AI (CeRAI), Wadhwani School of Data Science and AI (WSAI) Indian Institute of Technology Madras *krithishailya01@gmail.com, +gokul@cerai.in
1 Introduction
The integration of Large Language Models (LLMs) into educational guidance systems represents a paradigm shift in how students access academic advice. These systems promise access to personalized university and program recommendations, potentially addressing traditional barriers to quality educational counseling Ramos Pinho and Primo (2023), Chen et al. (2024). However, the deployment of LLMs in high-stakes educational decisions raises critical questions about fairness, representation, and the perpetuation of existing inequalities.
LLMs are trained on vast, uncurated internet corpora that embed societal biases and structural inequalities, so they risk reproducing and amplifying these distortions in their outputs Blodgett et al. (2020). Although bias in LLMs has been extensively studied across domains Cheng et al. (2025), its implications for educational recommendations remain largely unexplored. This is alarming because university choice profoundly shapes career trajectories and socioeconomic mobility Carnevale et al. (2015). In many developing countries, there is a widespread belief that foreign degrees confer superior quality and job prospects Haldorai et al. (2017). At the same time, educational technology firms are deploying AI-powered chatbots to guide admissions which can amplify existing disparities if based on biased LLMs. When an LLM repeatedly steers all users toward elite Western institutions, ignoring their geographic, economic, or cultural context, it misguides students and entrenches global hierarchies. The “black-box” nature of these models further compounds this, since users cannot assess the fairness of the advice they receive Yan et al. (2024).
To address this gap, we present three key contributions:
  • Academic Recommendation Queries: A comprehensive empirical study examining bias patterns in university recommendations across three popular open-source LLMs, analyzing 10,800 queries spanning 40 nationalities, 3 economic classes, and 3 genders.
  • Novel Evaluation Metrics: We present a novel evaluation framework that consists of two metrics – Demographic Representation Score (DRS) and Geographic Representation Score (GRS) which quantify the recommendation quality through dual lenses of demographic fit and geographic diversity respectively, providing a structured approach to assess fairness in the task of academic/university recommendation.
  • Evaluation & Analyses: Through the proposed evaluation framework, we present empirical evidence of significant biases across all evaluated models, with quantitative benchmarks that can guide future fairness research such as bias mitigation in LM based systems in educational sector. We make these query prompts public and evaluation framework as a benchmark.  [^1] https://github.com/cerai-iitm/Academic-Recommendation-Framework
Figure 1: Demographic and geographic biases in university and program recommendations. X and Y represent controlled demographic placeholders in this setup.
Our findings reveal interesting yet concerning patterns that could potentially impact the academic ecosystems across the globe both from a student and university/country perspective as shown in Figure 1. All models exhibit strong Western-centric bias, with 52–80% of recommendations favoring institutions in the United States (U.S.) and the United Kingdom (U.K.). Typical gender-stereotypical suggestions are prevalent – female profiles are steered toward social sciences and development studies, males toward engineering and computer science, and transgender users disproportionately to gender studies and social work. Economic status correlates with institutional prestige, potentially reinforcing socioeconomic barriers. These results highlight the urgent need to address bias and improve global representation in educational LMs.
2 Related Work
Recommender systems have become integral tools across various sectors, from e-commerce to education, yet they often inherit and amplify existing societal biases. For instance, employment recommenders steer gender-varying fictitious profiles toward lower-wage roles, smaller firms, and gendered language, an effect traced largely to content-based matching on gender inputs Zhang and Kuhn (2024). Färber et al. (2023) further offer a taxonomy that separates biases originating in human decisions from those introduced by algorithmic design, a distinction directly applicable to educational recommendation contexts.
Geographical bias similarly pervades AI. In relocation, tourism, and entrepreneurship prompts, LLMs systematically over- and under-represent certain locales, reinforcing a “rich-get-richer” effect Dudy et al. (2025). U.S. models perform up to 300% worse on salary, employer, and commute predictions in smaller metros than in the largest ones Campanella and Van Der Goot (2024). Globally, travel and story prompts mention poorer countries far less frequently and in more negative terms than wealthier ones Bhagat et al. (2024), mirroring the “US bias” observed in image generators Basu et al. (2023). Recent metrics comparing geographical and semantic distances reveal spatial distortions across ten major LMs Decoupes et al. (2024), and audits confirm under-representation of lower-socioeconomic regions Manvi et al. (2024).
Despite these insights, bias in educational recommendation systems remains under-studied. Most studies focus on knowledge queries, and treat demographic factors in isolation. Controlled, intersectional evaluations are needed to uncover how combined attributes like gender, class, and nationality shape LLM recommendations.
3 Methodology: Evaluation Framework for University Recommendations
Evaluating generative models in academic advising requires more than simple accuracy or relevance scores. A single metric can’t capture the complexity of a "good" recommendation, which must balance personalization, equity, diversity, and quality. To address this, we introduce a multi-dimensional evaluation framework that breaks recommendation quality into meaningful components, drawing from sociology, geography, and information retrieval.
The framework has two main pillars (Figure 2). Demographic Representation Score (DRS) measures how well recommendations fit a student’s background. Geographic Representation Score (GRS) evaluates overall set-level representation and quality among the global pool of universities. By examining each component, we gain detailed insights into a model’s behavior and biases.
Figure 2: Overview of the key perspectives and components of the evaluation framework
3.1 Quantifying Student-Centric Fit: Demographic Representation Score (DRS)
DRS measures how well a model can recommend universities that align with a prospective student’s profile consisting of demographics and academic details. It includes three metrics: Socio-Economic Accessibility ($Acc$), Reputation Alignment ($Rep$), and Academic Program Alignment ($Acad$).
3.1.1 Socio-Economic Accessibility
The Accessibility score models the socio-economic fit between a student $s$ and a university $u$ via:
$Acc(s,u)=e^{-\lambda\cdot d(s,u)}$ (1)
where $\lambda$ is a decay parameter and $d(s,u)$ is the geodesic distance (in km) between the capital cities of the student’s and university’s countries, calculated using Vincenty’s formula Vincenty (1975) via the geopy library, providing approximate structural distance between a student and institution.
This applies the distance-decay principle, an algorithm relating distance to utility Verma and Ukkusuri (2025). Here, we repurpose this concept to model the decay of educational opportunity over a socio-economic distance. Values near 1 denote perfect accessibility (zero distance), while values near 0 indicate extreme inaccessibility.
The decay parameter $\lambda$ acts as a socio-economic sensitivity controller. A larger $\lambda$ represents steeper barriers to accessfitting for low-income students, while a smaller $\lambda$ simulates scenarios with greater mobility. Based on our experiments, to ensure enough variance, we use $\lambda=0.0001$ for high class, $0.0005$ for middle class, and $0.001$ for low class profiles. This also allows our framework to reflect varied socio-economic realities and can be adapted to different national contexts.
3.1.2 Reputation Alignment
The Reputation Alignment score quantifies the institutional prestige of a recommended university based on established global or national ranking systems. It is calculated via linear normalization:
$Rep(u)=\frac{R_{max}-R_{u}}{R_{max}-R_{min}}$ (2)
where $R_{u}$ is the university $u$’s rank, and $R_{min}$ and $R_{max}$ are the best and worst ranks in the ranking system. Based on the scope of the QS rankings, we set a ceiling of $R_{max}$ as 1200 and $R_{min}$ as 1. Any university ranked beyond this threshold, or not ranked at all, receives a reputation score of 0.
This metric captures institutional quality and prestige, key factors in student choice and later outcomes Dale and Krueger (2002). The above formula converts raw rankings, where a lower number is better, into an intuitive score from 0 to 1 with a higher score indicating a better rank.
In conjunction with the $Acc$ score, $Rep$ reveals a model’s classification strategy: high $Rep$ but low $Acc$ (“Prestige-Seeking”) neglects student constraints, low $Rep$ but high $Acc$ (“Constraint-Adherent”) limits student aspirations; and a “Balanced” approach aligns prestige with accessibility.
3.1.3 Academic Alignment
The Academic Alignment score measures the curricular fit between a student’s interests and a university’s offerings. It is defined using a formula analogous to a Jaccard index Travieso et al. (2024).
$Acad(s,u)=\frac{|T_{s}\cap T_{u}|}{|T_{s}\cup T_{u}|}$ (3)
where $T_{s}$ is the set of subject tags for the student’s interests and $T_{u}$ is the set of subject tags for the university’s recommended programs.
The metric provides a measure for content-based relevance, ensuring that recommendations are not just prestigious or affordable but also aligned with the student’s academic goals. A score of 1 indicates a perfect match, while 0 indicates no overlap.
The complete DRS is formulated as a weighted arithmetic mean of its sub-metrics.
$DRS=w_{1}\cdot Acc+w_{2}\cdot Rep+w_{3}\cdot Acad$ (4)
where $w_{1}+w_{2}+w_{3}=1$ are the weights assigned to each component. While the framework allows for flexible weighting schemes to emphasize different aspects based on context (e.g., prioritizing accessibility for marginalized groups), in this work we adopt an equal weighting strategy.
However, for the purpose of model analysis, we also focus on the behavior of three individual components, as they reveal critical trade-offs in the recommendation task. Evaluating them in isolation lets us assess a model’s ability to balance aspiration and practicality, rewarding those that identify institutions both “aspirational” and “accessible”.
3.2 Assessing Geographic Diversity: Geographic Representation Score (GRS)
The GRS components evaluate the properties of the entire set of recommended universities. Their purpose is to assess how well the recommendation set represents the higher education landscape of a given country, enforcing a balance between the breadth of coverage and the reputational quality of the included institutions.
3.2.1 Sub-Metric: Normalised Representation
This metric is a ratio of two underlying components: Representation and Availability.
Representation ($Repr$) measures the proportion of a country’s ($c$) universities that were recommended by a model at least once.
$Repr(c)=\min\left(1.0,\frac{|Recs_{c}|}{|Total\_Unis_{c}|}\right)$ (5)
where $|Recs_{c}|$ is the number of recommended universities in country $c$, and $|Total\_Unis_{c}|$ is the total universities in our catalog for that country. This metric evaluates diversity by rewarding models that sample from a wider range of institutions.
Availability ($Avail$) establishes a baseline weight for each country, reflecting the relative size of its higher education sector.
$Avail(c)=\frac{|Total\_Unis_{c}|}{|Total\_Unis_{Global}|}$ (6)
where the denominator is the total number of universities across all countries in the QS rankings.
The final metric, the Normalised Representaion is defined as:
$Scaled\_Repr(c)=\min\left(1.0,\frac{Repr(c)}{Avail(c)+\epsilon}\right)$ (7)
where $\epsilon$ is a small constant (1e-6) to ensure numerical stability. A score greater than 1 (clipped to 1.0) indicates a country is being over-represented relative to its available set of universities, A score less than 1 indicates under-representation, despite having accessible options within the country.
This tackles a key source of bias in global recommender systems: the dominance of countries with large higher education sectors Yi et al. (2019). An LLM trained on web data will encounter vastly more text about U.S. universities than those in Brazil. Without normalization, a model would be rewarded for this biased recall. By adjusting for each country’s academic system size, we ensure fairer comparisons and test a model’s ability to draw on knowledge beyond training distributions.
3.2.2 Sub-Metric: Reputational Coverage
This metric acts as a qualitative guardrail, ensuring that a model’s representation of a country is not achieved by recommending only low-quality or obscure institutions.
$Rep\_covg(c)=\frac{\sum_{u\in Recs_{c}}\text{count}(u)\cdot Rep_{local}(u)}{\sum_{u\in Recs_{c}}\text{count}(u)}$ (8)
where $count(u)$ is the total number of times university u was recommended for country $c$, and $Rep_{local}(u)$ is its normalized reputation score as defined previously, but with the $R_{max}$ and $R_{min}$ as the max and min ranks of a particular country. This ensures that even if countries do not have high reputation universities overall, the model should be awarded for ranking the best universities in their coverage. This metric rewards models that not only name many universities within a country but also frequently recommend those of high repute.
A model could achieve a high $Repr$ score by suggesting three colleges, but if none of them are reputed, its $Rep_{covg}$ score would almost 0. To achieve high representation, a model should recommend less-common universities. To achieve high reputational coverage, it should stick to the well-known list. A model that balances these competing objectives will produce a recommendation set that is of recognized quality from diverse institutions.
The complete GRS is calculated as the geometric mean of its components, a choice that penalizes imbalance heavily, ensuring a high score cannot be achieved by excelling in one aspect while failing in another.
$GRS(c)=\sqrt{Scaled\_Repr(c)\cdot Rep\_covg(c)}$ (9)
4 Experimental Design
This section presents a reproducible experimental protocol showcasing our evaluation metrics’ utility. We detail constructing a global university knowledge corpus, generating synthetic user profiles and integrating them into our prompt templates. We then introduce these prompts on the target LLMs and their performance on our proposed academic metrics. We then outline our prompting strategy and the technical implementation details, including all hyperparameters used for generation.
4.1 University Knowledge Corpus
4.1.1 Institutional Data and Rankings
To create a comprehensive list of globally recognized institutions, we source university names, locations (country), and prestige rankings from the 2024 QS World University Rankings, specifically chosen to accurately test the model in accordance to the time of their release. A total of 1503 unique universities from over 120 countries were compiled.
4.1.2 Academic Program Data
We defined an Academic Alignment ($Acad$) score using a subject-tag taxonomy based on the five QS World University Rankings by Subject categories: Arts & Humanities, Engineering & Technology, Life Sciences & Medicine, Natural Sciences, and Social Sciences & Management. This provides a standardized and academically recognized classification scheme. Given the vast and inconsistent nomenclature of Master’s programs generated by the models (e.g., “MSc in Data Science”, “Master of Information and Data Science”), we prompted a large LLM (Llama-3-70B-Instruct) with a few annotated examples to assign one or more of these five tags to each program name. This method may be limited by potential biases in the auxiliary AI classifier, which we mitigate using manual review.
4.2 Synthetic User Profile Generation
To conduct a controlled experiment and isolate the impact of specific demographic attributes, we systematically generated a comprehensive set of synthetic user profiles. This approach avoids the ethical and privacy concerns of using real user data while enabling a thorough, intersectional analysis. Each profile was constructed by combining values from three demographic categories, as illustrated in Figure 3 and detailed below:
  • gender: The inclusion of a non-binary gender identity is critical for assessing the model’s inclusivity beyond traditional binaries.
  • economic_class: These terms serve as proxies for socioeconomic status (SES).
  • nationality: A diverse set of 40 nationalities was selected for global representation detailed in Appendix A.
The complete combination of these attributes resulted in 360 unique user profiles (3 Genders × 3 Economic Classes × 40 Nationalities).
4.3 Target Models
We evaluate three instruction-tuned, open-source LLMs, Llama-3.1 AI (2024), Gemma Team et al. (2024), and Mistral Jiang et al. (2023), chosen for their research popularity, open accessibility (vital for reproducibility), and diverse origins (Meta, Google, Mistral AI). Their similar size ( 7–8B parameters) lets us compare biases without scale confounds. We focus on smaller models both for computational efficiency and because lightweight LLMs are more practical for real-world chatbot deployments.
4.4 Prompting strategy
We designed three prompt templates, illustrated in Figure 3, to evaluate baseline biases and the impact of simple user-side interventions.
Table 1: Demographic Representation Score and its components for the base prompt, by demographic factors
Category Group Gemma Llama Mistral
Access. Rep. DRS Access. Rep. DRS Access. Rep. DRS
Overall Avg. 0.1336 0.5922 0.3629 0.3146 0.8479 0.3875 0.1786 0.7355 0.4570
Gender Male 0.1414 0.6058 0.3736 0.2967 0.8469 0.3812 0.1829 0.7310 0.4569
Gender Female 0.1728 0.5977 0.3652 0.3245 0.8390 0.3878 0.1965 0.7703 0.4834
Gender Transgender 0.1267 0.5732 0.3499 0.3227 0.8577 0.3935 0.1563 0.7052 0.4307
Economic Class High-Class 0.1252 0.6543 0.3897 0.2651 0.9044 0.3898 0.1500 0.9638 0.5569
Economic Class Moderate-Class 0.1318 0.5711 0.3515 0.3225 0.8460 0.3895 0.1897 0.6701 0.4299
Economic Class Low-Class 0.1439 0.5513 0.3476 0.3563 0.7932 0.3832 0.1960 0.5726 0.3843
Figure 3: Prompt template and augmentation setup used for university recommendation experiments
The base template is a standard university recommendation query with demographic placeholders. The regional accessibility augmentation adds an explicit constraint to counter Western-centric bias and test model’s ability to adapt to user’s geographic context. The educational background augmentation tests for recommendations aligned with the user’s skills. We also conducted a reduced-context experiment, providing only a single demographic attribute, as detailed in Appendix B.
For each of the 360 unique user profiles, every prompt template was queried on three separate models. To account for stochasticity and ensure fair comparison, each prompt-model pair was run 10 times using identical decoding parameters and strict output formatting instructions.
4.5 Implementation
All experiments are conducted with defined parameters to ensure reproducibility. We use a temperature of 0.75 and run our evaluation in Python 3.10, loading models from the hugging face transformers library. The detailed setup is given in Appendix A.
5 Results and Discussion
For each prompt, we generate a list of three universities from the target LLMs as academic recommendations on which we compute the proposed suite of disaggregated metrics.
For each of the three recommended universities, we calculate $Acc$, RepRep, and $Acad$ and report the average of these scores across the recommendations. We also calculate the set-level metrics $Repr/Avail$ and $Rep_{cov}$ for the specified target country.
We analyzed the results and discuss them under the following Research Questions (RQ). Further results in detail are shown in Appendix B.
5.1 RQ1: Do LLM Recommendations Reflect and Reinforce patterns based on Demographic and Economic Status?
Our findings highlight that LLMs are far from being neutral information arbiters, and act as mirrors that reflect and amplify societal stereotypes about class and gender. This is starkly evident in their creation of distinct recommendation "tiers" based on a user’s perceived socio-economic status.
5.1.1 Economic Class
Table 1 exposes clear socio-economic stratification where models prioritize prestige over practicality for "High-Class" profiles, models, with Mistral recommending universities with a high Reputation score but low Accessibility. For "Low-Class" profiles, Mistral’s recommendations invert, with Reputation plummeting by 41%, which also holds across models. Llama’s score for high-class profiles is 1.14 times higher than for low-class. This amounts to ‘digital gatekeeping’: models preemptively filter out top-tier options to lower-income backgrounds, despite numerous scholarships opportunities offered by institutes, filtering opportunities based on a demographic proxy, rather than merit.
Table 2: Detailed Geographic Representation Score (GRS) for select countries, grouped by development status
Gemma Llama Mistral
Country Avail. Repr. Rep. Covg. GRS Repr. Rep. Cov. GRS Repr. Rep. Cov. GRS
Developed Nations
   Canada 0.0200 0.2333 0.9698 0.9848 0.7000 0.9347 0.9668 0.5667 0.9189 0.9586
   United Kingdom 0.0599 0.2444 0.9882 0.9941 0.8222 0.8992 0.9483 0.5333 0.8994 0.9484
   United States 0.1311 0.1066 0.9731 0.8896 0.2386 0.9253 0.9619 0.4315 0.9123 0.9552
Developing Nations
   South Africa 0.0073 0.3636 0.8413 0.9172 1.0000 0.7022 0.8379 0.5455 0.7443 0.8627
   Nigeria 0.0013 0.0000 0.0000 0.0000 1.0000 0.0829 0.2880 0.0000 0.0000 0.0000
   India 0.0306 0.0000 0.0000 0.0000 0.0217 0.0000 0.0000 0.0000 0.0000 0.0000
Table 3: Comparison of Academic Alignment Scores Across Demographic Groups and Models
Group Gemma Llama Mistral
By Gender
   Female 0.4451 0.6866 0.7174
   Male 0.5127 0.7851 0.7903
   Transgender 0.3539 0.6257 0.7242
By Economic Class
   High-Class 0.4506 0.6334 0.7206
   Moderate-Class 0.4228 0.6729 0.7147
   Low-Class 0.4183 0.5912 0.6906
5.1.2 Gender
This trend extends to gender, where quantitative metrics reveal damaging biases. As shown in Table 3, academic alignment shows a consistent disparity. Both Llama and Gemma provide male profiles with recommendations better aligned to their interests than female profiles. The gap is most alarming for transgender users, where Gemma’s score plummets to 0.3539. This numerical gap represents a tangible failure, detailed in Figure 4: a transgender user asking for "Computer Science" is more likely to be recommended misaligned programs like "Social Work," rendering the advice functionally useless. Recommendations adhere to rigid gender stereotypes, steering men towards engineering while funneling women and transgender profiles into social policy. This bias persists even when a prompt emphasizes a strong engineering background; women and transgender users still receive many social-policy suggestions. This persistence, resistant to simple alignment, shows how gender and geography distort the model’s advice. Ultimately, the model’s stereotypical associations override the user’s defined skillset defeating the fundamental purpose of a recommendation system.
Figure 4: Program Recommendation trends by gender under the base prompt and the context-augmented prompt (A) with engineering background.
Table 4: Impact of the ’Regional’ Prompt on GRS, for a select few nations.
Gemma Llama Mistral
Country Base Regional $\Delta$ (%) Base Regional $\Delta$ (%) Base Regional $\Delta$ (%)
Developed Nations
Canada 0.9848 0.9895 +0.5% 0.9668 0.9895 +2.4% 0.9586 0.9895 +3.2%
Australia 0.0000 0.9946 $\infty$ 0.9457 0.7733 -18.2% 0.9517 0.9921 +4.2%
Italy 0.8972 0.0000 -100% 0.8103 0.0000 -100% 0.0000 0.0000 0%
Japan 0.9713 0.0000 -100% 0.8644 0.0000 -100% 0.0000 0.2880 $\infty$
Germany 0.0000 0.0000 0% 0.9178 0.0000 -100% 0.9767 0.0000 -100%
Developing Nations
Ghana 0.0000 0.5204 New 0.5204 0.5783 +11.1% 0.5204 0.5328 +2.38%
Nigeria 0.0000 0.3400 New 0.2880 0.3800 New 0.0000 0.3720 New
South Africa 0.9172 0.0000 -100% 0.8379 0.0000 -100% 0.8627 0.0000 -100%
Philippines 0.8485 0.0000 -100% 0.6801 0.0000 -100% 0.0000 0.0000 0%
India 0.0000 0.0000 0% 0.0000 0.0000 0% 0.0000 0.0000 0%
Brazil 0.0000 0.0000 0% 0.0000 0.0000 0% 0.0000 0.0000 0%
5.2 RQ2: Can I trust an LLM to give me recommendations that are representative of the global education sphere?
The LLMs’ recommendation base is a profoundly incomplete and distorted world map, leaving vast regions in a representational shadow. The most representative model, Llama-3.1-8B, covers less than half the globe (48%), while Gemma’s worldview is a meager 17.4% of countries, severely limiting the scope of possible recommendations.
The consequences of this distorted cartography are quantified by the Geographic Representation Score (GRS) in Table 2 and qualitatively detailed in Appendix B. A small cohort of Western nations constitutes the models’ "known world," receiving high GRS scores and excellent Reputational Coverage (often > 0.90), signifying that the models can name a diverse and high-quality set of institutions within these countries. In contrast, most of the world is a blank space. For nearly all developing nations testes, Gemma and Mistral return a GRS of zero. Countries like India, despite a massive higher education system, are rendered completely invisible with a GRS of zero across all models.
Even when a model appears aware of the Global South, the sub-metrics highlight that this is dangerously superficial. Llama gives Nigeria a perfect Representation (Rep) of 1 but a weak Reputation Coverage of only 0.0829. The model can name a university, but not reliably a good one, offering users a harmful illusion of competence.
Table 5: Comparison of DRS and sub-metrics for Base (B) and Regional (R) prompts across models.
Model (Prompt) DRS Acc Rep
Gemma (B) 0.3664 0.1336 0.5922
Gemma (R) 0.2252 0.1493 0.3011
Llama (B) 0.5812 0.3146 0.8479
Llama (R) 0.4707 0.3969 0.5446
Mistral (B) 0.4570 0.1786 0.7355
Mistral (R) 0.3316 0.1669 0.4963
5.3 RQ3: Can User-Side Prompt Engineering Overcome Systemic Representational and Stereotypical Deficits?
Our setup also introduces a “regionally-accessible” constraint to test if user-side prompt engineering could mitigate systemic flaws. The results (Tables 4 and 5) show this is not a simple fix and can yield unpredictable, even detrimental, outcomes.
Across all models, adding the regional prompt decreased the overall DRS because the significant drop in university Reputation outweighed modest gains in Accessibility. While this was expected, models constrained geographically fell back on lesser prestigious institutions than from previously recommended regions, thus lowering the quality, visible in some nation trends like South Africa and the Philippines reduced to null scores.
Some previously underrepresented nations like Nigeria gain visibility and Australia gains more reputed universities resulting in a higher GRS. This demonstrates that for some regions, the models have a degree of latent knowledge that needs explicit direction which is also highly unstable. For Llama, representation for major developed nations like Italy, Japan, and Germany (with strong base GRS scores >0.81) collapsed entirely to 0.0000.
Crucially, major developing nations like India and Brazil still scored GRS=0 across all models, even under regional constraints. Likewise, adding academic context failed to overcome biases, confirming that user-side prompts alone cannot bridge these knowledge gaps. Our framework thus also points towards bias mitigation strategies like fairness-aware losses or essential context data required. While tested for higher-education recommendations, our socially grounded framework can be applied to other tasks like in Appendix C where accessibility and reputation are vital aspects in recommendation systems.
6 Conclusion
This paper delivers a comprehensive analysis of how open-source LLMs shape higher-education recommendations, uncovering stark demographic and geographic biases. By applying our replicable framwork with Demographic and Geographic Representation Scores, we quantify unfairness showing that models favor high-class users with prestigious yet inaccessible universities and filter out low-class profiles; gender misalignment persists despite tailored prompts, and major education hubs like India and Brazil remain invisible. Among the tested models, Llama is the most globally representative, while Gemma performs worst. This work presents an instrumental step towards building equitable academic AI that ensures that every student, regardless of background, receives recommendations that are both aspirational and attainable.
7 Limitations
While this analysis represents a comprehensive examination of bias in LLM-based academic recommendation, there are a few limitations to be considered:
  • Synthetic Profile Scope. Our 360 synthetic profiles enable controlled, intersectional analysis across gender, nationality, and socioeconomic status, but cannot capture real-world complexity such as scholarships, dual-degree plans, or personal constraints.
  • Dependence on QS Rankings. Both Reputation and Geographic Representation metrics rely on the 2024 QS World University Rankings; any omissions or biases in that dataset, particularly undercoverage of emerging universities, directly affect our results.
  • Subject-Tag Taxonomy Reliance. Program titles are mapped into five broad QS subject areas via a secondary LLM and manual checks. This standardization brings consistency but can introduce noise, especially for interdisciplinary or novel programs, slightly affecting Academic Alignment scores.
  • Model and Scale Constraints. We evaluate three 7–8 B-parameter open-source LLMs; findings may not extend to larger foundation models (30 B+), closed-source systems (e.g., GPT-4 & Gemini), or domain-tuned variants, which may exhibit different biases.
  • Fixed Decay Parameters. The decay constants $\lambda$ for high, moderate, and low economic classes were chosen to generate variance in Accessibility scores but remain heuristic and may not reflect real financial or visa barriers.
  • Unmeasured Intersectional Axes. We vary gender, nationality, and economic status, but other factors like language proficiency, disability also shape educational opportunity which needs further research and can be included in future work.
8 Ethical Considerations
Our evaluation framework goes a step further than standard metrics by providing different perspectives for practitioners to understand what a model lacks. The integration of our Demographic Representation Score (DRS) and Geographic Representation Score (GRS) into LLM-based recommendation systems reflects a commitment to understanding and mitigating the real-world impacts of algorithmic advice. Unlike traditional evaluation metrics that focus solely on accuracy or relevance, DRS and GRS illuminate how well model outputs align with students’ socioeconomic constraints, personal interests, and the full breadth of global higher education.
Since these metrics are calculated group level, they also help analyse what country/attribute is under represented to level down and analyse a model’s strengths and weaknesses to train them with the right type of data. In practice, a high DRS score signals to developers that their system is successfully tailoring suggestions to a student’s unique context, rather than defaulting to one-size-fits-all “elite” or “popular” choices. Conversely, a low DRS immediately highlights demographic blind spots, such as systematic exclusion of lower-income profiles or misalignment with expressed program interests, prompting targeted data curation or re-weighting of loss functions.
Similarly, GRS goes beyond mere country counts by normalizing representation against each nation’s landscape of accredited universities. A model with a robust GRS does not merely recall a handful of well-known Global North institutions, it surfaces a diverse mix of universities that collectively reflect regional availability and quality. Institutions can use GRS to audit their own AI-driven advising tools, ensuring that education systems are equally represented and receive fair consideration. Policymakers and accreditation bodies may likewise reference GRS benchmarks when certifying digital counseling platforms, embedding fairness metrics into compliance standards.
Our framework is designed for broad applicability. University career centers and online counseling platforms can adopt DRS and GRS as part of their continuous integration pipelines, comparing new model versions against fairness baselines before deployment. It also helps users decide what models are best setting a new standard for fairness evaluation in educational recommendation contexts.
Beyond higher education, the principles underlying DRS and GRS extend naturally to other recommendation domains, job matching services, healthcare provider selection, or financial product advisories, where balancing user constraints, domain expertise, and population-level diversity is equally critical. Given an official-sourced ranking data, this social taxonomoy can also be extended to these domains to evaluate similar representational and demographic bias detailed in Appendix C.
By embedding DRS and GRS into the development lifecycle of educational recommendation systems and by articulating their intended uses, limitations, and potential pitfalls, we foster a more transparent, accountable, and equitable ecosystem for AI-driven guidance. Our work strongly highlights the urgent need to overcome systemic knowledge deficits through deeper methods like algorithmic de-biasing, curriculum-aware fine-tuning, and enriched non-Western training corpora. Through open release of code and data splits and collaborative refinement of these metrics will be essential to ensure that algorithmic advising genuinely advances access to quality education for all.
Acknowledgements
We are grateful to Mr. Krishnan Narayanan (Ihihaasa/IIT Madras), Prof. Vanessa Evers (University of Twente, Netherlands), Dr. Akuadasuo Ezenyilimba (Arizona State University, USA), Dr. Gerlinde Kristahn, Global Learning Council (Villars-sur-Ollon, Switzerland) and our fellow researchers from Centre for Responsible AI at IIT Madras for their valuable suggestions and critical inputs.
Appendix A Experimental Setup
A.1 Models
In this study, we evaluated three prominent open-source, instruction-tuned Large Language Models (LLMs). The models were selected based on their wide adoption in the research community, open accessibility which is crucial for reproducibility, and their diverse origins, allowing for a comparative analysis. Their similar scale ( 7-8B parameters) ensures that our comparisons of bias are not confounded by model size. We focused on these smaller models due to their computational efficiency and practical relevance for real-world chatbot deployments.
The specific models used are:
  • Llama-3.1-8B-Instruct: Created by Meta (version released July 23, 2024). Accessed via the Hugging Face Hub at meta-llama/Meta-Llama-3.1-8B-Instruct.
  • gemma-7b-it: Created by Google. Accessed via the Hugging Face Hub at google/gemma-7b-it.
  • Mistral-7B-Instruct-v0.3: Created by Mistral AI. Accessed via the Hugging Face Hub at mistralai/Mistral-7B-Instruct-v0.3.
Our use of these models is fully consistent with their intended use for research and experimentation. The evaluation of model biases and limitations aligns with the responsible AI development practices encouraged by their creators. These instruction-tuned models are designed for a wide range of natural language generation tasks. Our study uses them in a research context to evaluate their performance, biases, and alignment capabilities on a specific, high-stake task (academic advising). This falls squarely within the intended scope of research and experimentation encouraged by the model creators. Our usage complies with the Acceptable Use Policies of both Llama 3.1 and Gemma, as our experiments do not involve any prohibited activities such as generating illegal content, hate speech, or misinformation. The purpose of our work is to identify and analyze potential harms (i.e., bias), which is a crucial aspect of responsible AI research. To ensure the privacy and ethical integrity of our study, we avoided using any real user data.
The models are governed by distinct open licenses that permit research use: Llama-3.1-8B-Instruct is licensed under the Llama 3.1 Community License Agreement, Gemma-7b-it is governed by the Gemma Terms of Use, and Mistral-7B-Instruct-v0.3 is released under the permissive Apache 2.0 License. Our use of these models is fully consistent with their intended use for research and experimentation. The evaluation of model biases and limitations aligns with the responsible AI development practices encouraged by their creators and complies with the Llama 3.1 Acceptable Use Policy and the Gemma Prohibited Use Policy.
A.2 Computing Requirements
The experimental pipeline was implemented in Python 3.10. Models were loaded and queried using the Hugging Face transformers library (v4.38.2) with the PyTorch (v2.1) backend. All experiments were executed on the Kaggle, utilizing notebooks equipped with NVIDIA T4 GPUs to accelerate inference. Data processing and analysis were conducted using the pandas and numpy libraries.
A total of 32,400 model generations were performed (360 profiles × 3 prompts × 3 models × 10 runs). The total computational budget is estimated to be approximately 45-50 GPU hours on the specified hardware.
This study evaluates pre-trained models, so no model training or fine-tuning was performed. The key hyperparameters relate to the text generation (decoding) process. To ensure a fair and consistent comparison across all models, a fixed set of decoding parameters was used for every query detailed in Table 6.
To account for the stochastic nature of generative models, each unique prompt-model configuration was queried 10 independent times. This approach provides a stable and representative measure of each model’s typical behavior, mitigating the randomness inherent in a single generation. While not included in the tables for brevity, this multi-run setup allows for the calculation of variance and standard deviation around the reported means.
To ensure reproducibility, specific versions of all major software packages were used. No modifications were made to the core functionalities of these libraries.
  • Core ML/DL Libraries: transformers (v4.38.2), torch (v2.1).
  • Data Handling: pandas (v2.0.3), numpy (v1.25.2).
  • Geospatial Calculations: geopy (v2.4.1) was used to calculate the geodesic distance for the Socio-Economic Accessibility (Acc) score.
Parameter Value
temperature 0.75
top_p 0.95
max_new_tokens 300
do_sample True
num_return_sequences 1
Table 6: Decoding hyperparameters used for all model queries.
A.3 Prompt Details
Countries used in prompt template: Africa, Asia, Europe, North America, South America, and Oceania. The list includes: Nigeria, Egypt, South Africa, Kenya, Ghana, Ethiopia, Algeria, Morocco, China, India, Japan, South Korea, Indonesia, Thailand, Saudi Arabia, Vietnam, France, Germany, Italy, Spain, United Kingdom, Sweden, Poland, Greece, United States, Canada, Mexico, Cuba, Costa Rica, Jamaica, Brazil, Argentina, Chile, Peru, Colombia, Australia, New Zealand, Fiji, Papua New Guinea, and Tonga.
Appendix B Qualitative Analysis
This section lays the qualitative analysis of the models’ performance on different prompt variations based on the demographic factors like gender, economic-class and nationality of a simulated student seeking academic advice.
B.1 Base Prompt
The volume of data generated from the base prompt is tabulated in Table 7:
Table 7: Volume and diversity of generated responses for the base prompt template.
Gemma 7B LLaMA 3.1 8B Mistral 7B
Total Responses 6,900 13,176 10,994
Unique Universities 96 481 229
Unique Programs 296 1309 814
Unique Countries 22 61 27
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 5: Distribution of the top 20 most frequently recommended university locations across the three models (Mistral, LLaMA, and Gemma).
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 6: The geographic spread of all universities recommended by the Mistral, LLaMA, and Gemma models reveals a strong Western bias, with a predominant focus on institutions from the United States, United Kingdom, Canada, and Australia.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 7: The count plot shows the top 20 universities most commonly suggested overall by the Mistral, LLaMA, and Gemma models.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 8: The most frequently recommended universities for each financial class by the Mistral, LLaMA, and Gemma models, revealing a strong influence of economic class in their recommendations.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 9: Frequency plot showing the top 20 academic programs recommended overall across the three models: Mistral, LLaMA, and Gemma.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 10: Academic program recommendations are grouped by gender for the Mistral, LLaMA, and Gemma models. Transgender users face the strongest bias across all three models.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 11: The most commonly recommended programs by economic status are shown for the Mistral, LLaMA, and Gemma models. The results indicate that program recommendations vary notably by users’ socioeconomic background.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 12: The heatmap shows the alignment between users’ nationality and the locations of recommended universities for selected nationalities. The models tend to favor institutions in developed countries, reflecting a Western-centric bias that underrepresents universities from the Global South.
B.2 Added Context of Regional Accessibility
The volume of data generated from the prompt with an additional context of regional accessibility is tabulated in Table 8:
Table 8: Volume and diversity of generated responses for the prompt with additional regional context.
Gemma 7B LLaMA 3.1 8B Mistral 7B
Total Responses 6,077 26,794 9,623
Unique Universities 129 382 257
Unique Programs 127 423 245
Unique Countries 37 60 43
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 13: The frequency distribution of the top 20 recommended universities by location across the Mistral, LLaMA, and Gemma models, with additional regional context provided.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 14: Global spread of universities recommended by the Mistral, LLaMA, and Gemma models with consideration of regional accessibility. The models predominantly favor Western institutions, reflecting existing global academic hierarchies.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 15: The top 20 universities most commonly suggested overall by the three models when users request regional options. Despite the prompt, all models continue to prioritize prestigious Western institutions.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 16: Most frequently recommended universities for each financial group for the Mistral, LLaMA, and Gemma models, with accessibility taken into account. While some regional improvements are observed, all models align recommendations with income level, reinforcing educational inequality.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 17: The top 20 recommended programs, constrained by regional accessibility, highlighting persistent disciplinary biases across the Mistral, LLaMA, and Gemma models.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 18: Top program recommendations by gender identity across the Mistral, LLaMA, and Gemma models, with additional regional context, revealing systemic bias, with transgender users consistently steered toward stereotyped disciplines.
Figure 19 shows the comparative performance of two prompts(with and without regional context). Contextual prompts reduce Western bias in model recommendations, yet some countries remain underrepresented.
(a) Mistral
(b) LLaMA
(c) Gemma
Figure 19: Comparison chart of user’s nationality and university location alignment, with and without the regional accessibility cue in the prompt (selected nationalities).
B.3 Results for Prompt Template with Reduced Context (Individual Demographic Factors)
The following section presents the results obtained from the three models when the prompt template included only a single attribute at a time (i.e., either gender, economic class, or nationality). The outcomes are summarized in the tables for each model.
B.3.1 Results for Gemma-7B
The results for the Gemma-7B model, when prompted with templates containing only a single attribute (i.e., economic class, gender, or nationality), are presented in tables 9, 10, and 11.
Table 9: Prompted with only economic-class in the prompt: Gemma-7B.
Class Top Countries Top Universities Top Programs
Overall United States University of Oxford Public Policy
United Kingdom University of Chicago Economics
New Zealand UC Berkeley Business Administration
Low-class United States Boston University Public Policy
Moderate-class United States University of Chicago Public Policy
High-class United Kingdom University of Oxford Economics
Table 10: Prompted with only gender in the prompt: Gemma-7B.
Gender Top Countries Top Universities Top Programs
Overall United States Boston University Gender Studies
United Kingdom University of Oxford Public Policy
New Zealand Auckland University of Technology Business Administration in Economics
Male United States Boston University Business Administration in Economics
Female United States Boston University Business Administration
Trans United States UC Berkeley Gender Studies
Table 11: Prompted with only nationality in the prompt: Gemma-7B.
Nationality Top Countries Top Universities Top Programs
Overall United Kingdom University of Oxford Social Policy
United States University of Cambridge Development Studies
Australia University of East London Public Policy
US United States University of Chicago Business Administration
UK United Kingdom University of Oxford Social Policy
China United Kingdom University of Oxford Public Policy
Nigeria United Kingdom University of Oxford Social Policy
India United Kingdom University of Oxford Social Policy
Cuba United Kingdom University of Oxford Electrical Engineering
B.3.2 Results for LlaMA-3.1-8B
Tables 12, 13, and 14 present the results obtained from the LlaMA-3.1-8B model when prompted with templates that include only one attribute at a time.
Table 12: Prompted with only economic-class in the prompt: LlaMA-3.1-8B.
Class Top Countries Top Universities Top Programs
Overall United Kingdom University of Edinburgh Finance
Netherlands University of Oxford Data Science
United States University of Cambridge Economics
Low-class United Kingdom University of Edinburgh Data Science
Moderate-class United Kingdom University of Edinburgh Data Science
High-class United Kingdom University of Oxford Finance
Table 13: Prompted with only gender in the prompt: LlaMA-3.1-8B.
Gender Top Countries Top Universities Top Programs
Overall United Kingdom University of Edinburgh Data Science
United States University of Oxford Computer Science
Canada University of Cambridge Artificial Intelligence
Male United Kingdom University of Cambridge Computer Science
Female United Kingdom University of Edinburgh Data Science
Trans United Kingdom University of Edinburgh Gender Studies
Table 14: Prompted with only nationality in the prompt: LlaMA-3.1-8B.
Nationality Top Countries Top Universities Top Programs
Overall United Kingdom University of Edinburgh Data Science
United States University of Oxford Environmental Science
Australia University of Manchester Computer Science
US United Kingdom University of Edinburgh Data Science
UK United Kingdom University of Edinburgh Data Science
China United Kingdom University of Edinburgh Computer Science
Nigeria United Kingdom University of Edinburgh Data Science
India United Kingdom University of Edinburgh Computer Science
Cuba United Kingdom University of Edinburgh International Relations
B.3.3 Results for Mistral-7B
The outcomes generated by the Mistral-7B model in response to prompts containing a single attribute (economic class, gender, or nationality) are summarized in tables 15, 16, and 17.
Table 15: Prompted with only economic-class in the prompt: Mistral-7B.
Class Top Countries Top Universities Top Programs
Overall United States Stanford University Computer Science
United Kingdom Massachusetts Institute of Technology Data Science
UC Los Angeles Engineering Management
Low-class United States University of Texas at Austin Computer Science
Moderate-class United States UC Los Angeles Computer Science
High-class United States Stanford University Engineering Management
Table 16: Prompted with only gender in the prompt: Mistral-7B.
Gender Top Countries Top Universities Top Programs
Overall United States Massachusetts Institute of Technology Computer Science
United Kingdom UC Berkeley Data Science
Stanford University Social Work
Male United States Massachusetts Institute of Technology Computer Science
Female United States Massachusetts Institute of Technology Computer Science
Trans United States University of Michigan Ann Arbor Social Work
Table 17: Prompted with only nationality in the prompt: Mistral-7B.
Nationality Top Countries Top Universities Top Programs
Overall United States UC Berkeley Computer Science
United Kingdom University of Oxford Data Science
New Zealand Massachusetts Institute of Technology Artificial Intelligence
US United States UC Berkeley Computer Science
UK United Kingdom Imperial College London Computer Science
China United States Massachusetts Institute of Technology Computer Science
Nigeria United Kingdom University of Manchester Computer Science
India United States University of Illinois Urbana-Champaign Computer Science
Cuba United States UC Berkeley Computer Science
Appendix C Broader Application of Framework
The core principles of our evaluation framework, balancing accessibility, reputation, alignment, and diversity, are not limited to higher education. The social taxonomy introduced can be adapted to other high-stake recommendation domains where user context and equitable representation are critical. Below, we outline how the Demographic Representation Score (DRS) and Geographic Representation Score (GRS) can be re-conceptualized for other applications.
C.1 Job Recommendation Systems
For a job seeker, a "good" recommendation must balance commute, company quality, and skill match.
DRS Adaptation:
Socio-Economic Accessibility (Acc): This could be modeled as a function of the physical commute distance from the user’s home to the job location, or as a binary score for remote vs. in-person roles. The decay parameter $\lambda$ could represent a user’s willingness to commute.
Reputation Alignment (Rep): Instead of university rankings, this would use normalized company ratings from platforms like Glassdoor, or it could be based on publicly available salary-band data to represent economic opportunity.
Academic Alignment (Acad): Re-framed as Skill Alignment, this would use a Jaccard index to measure the overlap between a user’s skills (parsed from a CV) and the skills listed in the job description.
GRS Adaptation:
This score would evaluate the diversity of employers within a specific labor market (e.g., a city or region).
Normalized Representation (Scaled_Repr) would measure if a model recommends jobs from a wide range of companies relative to the total number of employers in that area, preventing over-concentration on a few large tech firms.
Reputational Coverage (Rep_covg) would ensure that the recommended companies are of high quality, based on the Rep score defined above.
C.2 Healthcare Provider Selection
Choosing a doctor or hospital involves balancing travel, quality of care, and specialty match.
DRS Adaptation:
Socio-Economic Accessibility (Acc): This could be a function of travel time to the clinic or hospital. More critically, it could also incorporate whether the provider is in the user’s insurance network, a crucial real-world accessibility barrier.
Reputation Alignment (Rep): This would be based on normalized patient satisfaction scores, official hospital safety grades, or professional accreditations from medical bodies.
Academic Alignment (Acad): Re-framed as Specialty Alignment, this would measure the match between a patient’s stated medical needs (e.g., "pediatric care," "cardiology") and the provider’s listed specialties.
GRS Adaptation:
This score would assess the diversity of recommended healthcare options within a health district or city.
Normalized Representation (Scaled_Repr) would check if the recommendations include a mix of large hospitals, specialized clinics, and local primary care physicians, relative to what is available.
Reputational Coverage (Rep_covg) would ensure that the recommended providers meet a high standard of care based on patient ratings or official grades.

### v2

@@
 Our findings highlight the urgent need for bias consideration in educational LMs to ensure equitable global access to higher education.
+We make these query prompts public and evaluation framework as a benchmark.
+[^1] https://github.com/cerai-iitm/Academic-Recommendation-Framework
 Where Should I Study? Biased Language Models Decide!
@@
 The integration of Large Language Models (LLMs) into educational guidance systems represents a paradigm shift in how students access academic advice.
-These systems promise access to personalized university and program recommendations, potentially addressing traditional barriers to quality educational counseling Ramos Pinho and Primo (2023), Chen et al.
-(2024).
+These systems promise access to personalized university and program recommendations, potentially addressing traditional barriers to quality educational counseling 10343248, chen_facilitating_2024.
 However, the deployment of LLMs in high-stakes educational decisions raises critical questions about fairness, representation, and the perpetuation of existing inequalities.
-LLMs are trained on vast, uncurated internet corpora that embed societal biases and structural inequalities, so they risk reproducing and amplifying these distortions in their outputs Blodgett et al.
-(2020).
-Although bias in LLMs has been extensively studied across domains Cheng et al.
-(2025), its implications for educational recommendations remain largely unexplored.
-This is alarming because university choice profoundly shapes career trajectories and socioeconomic mobility Carnevale et al.
-(2015).
-In many developing countries, there is a widespread belief that foreign degrees confer superior quality and job prospects Haldorai et al.
-(2017).
+LLMs are trained on vast, uncurated internet corpora that embed societal biases and structural inequalities, so they risk reproducing and amplifying these distortions in their outputs blodgett2020language.
+Although bias in LLMs has been extensively studied across domains cheng2025neutralizing, its implications for educational recommendations remain largely unexplored.
+This is alarming because university choice profoundly shapes career trajectories and socioeconomic mobility carnevale2015economic.
+In many developing countries, there is a widespread belief that foreign degrees confer superior quality and job prospects haldorai_determinants_2017.
 At the same time, educational technology firms are deploying AI-powered chatbots to guide admissions which can amplify existing disparities if based on biased LLMs.
 When an LLM repeatedly steers all users toward elite Western institutions, ignoring their geographic, economic, or cultural context, it misguides students and entrenches global hierarchies.
-The “black-box” nature of these models further compounds this, since users cannot assess the fairness of the advice they receive Yan et al.
-(2024).
+The “black-box” nature of these models further compounds this, since users cannot assess the fairness of the advice they receive yan2024practical.
 To address this gap, we present three key contributions:
@@
   • Evaluation & Analyses: Through the proposed evaluation framework, we present empirical evidence of significant biases across all evaluated models, with quantitative benchmarks that can guide future fairness research such as bias mitigation in LM based systems in educational sector.
-We make these query prompts public and evaluation framework as a benchmark.
-[^1] https://github.com/cerai-iitm/Academic-Recommendation-Framework
 Figure 1: Demographic and geographic biases in university and program recommendations. X and Y represent controlled demographic placeholders in this setup.
@@
 Recommender systems have become integral tools across various sectors, from e-commerce to education, yet they often inherit and amplify existing societal biases.
-For instance, employment recommenders steer gender-varying fictitious profiles toward lower-wage roles, smaller firms, and gendered language, an effect traced largely to content-based matching on gender inputs Zhang and Kuhn (2024).
-Färber et al.
-(2023) further offer a taxonomy that separates biases originating in human decisions from those introduced by algorithmic design, a distinction directly applicable to educational recommendation contexts.
+For instance, employment recommenders steer gender‑varying fictitious profiles toward lower‑wage roles, smaller firms, and gendered language, an effect traced largely to content‑based matching on gender inputs zhang_measuring_2024.
+farber2023biases further offer a taxonomy that separates biases originating in human decisions from those introduced by algorithmic design, a distinction directly applicable to educational recommendation contexts.
 Geographical bias similarly pervades AI.
-In relocation, tourism, and entrepreneurship prompts, LLMs systematically over- and under-represent certain locales, reinforcing a “rich-get-richer” effect Dudy et al.
-(2025).
+In relocation, tourism, and entrepreneurship prompts, LLMs systematically over‑ and under‑represent certain locales, reinforcing a “rich‑get‑richer” effect dudy2025unequal.
 U.S.
-models perform up to 300% worse on salary, employer, and commute predictions in smaller metros than in the largest ones Campanella and Van Der Goot (2024).
-Globally, travel and story prompts mention poorer countries far less frequently and in more negative terms than wealthier ones Bhagat et al.
-(2024), mirroring the “US bias” observed in image generators Basu et al.
-(2023).
-Recent metrics comparing geographical and semantic distances reveal spatial distortions across ten major LMs Decoupes et al.
-(2024), and audits confirm under-representation of lower-socioeconomic regions Manvi et al.
-(2024).
-Despite these insights, bias in educational recommendation systems remains under-studied.
-Most studies focus on knowledge queries, and treat demographic factors in isolation.
-Controlled, intersectional evaluations are needed to uncover how combined attributes like gender, class, and nationality shape LLM recommendations.
+models perform up to 300% worse on salary, employer, and commute predictions in smaller metros than in the largest ones campanella2024big.
+Globally, travel and story prompts mention poorer countries far less frequently and in more negative terms than wealthier ones bhagat2024richer, mirroring the “US bias” observed in image generators basu_inspecting_2023.
+Recent metrics comparing geographical and semantic distances reveal spatial distortions across ten major LMs decoupes2024evaluation, and audits confirm under‑representation of lower‑socioeconomic regions manvi2024large.
+Despite the growing body of research on bias in AI systems, several significant gaps remain in understanding geographical bias in educational recommendations specifically.
+Most existing studies on recommendation systems focus on e-commerce, job matching, or general information retrieval rather than educational contexts.
+Additionally, while geographical bias has been studied in various contexts, the intersection of geography with other demographic factors (gender, socioeconomic status, nationality) in educational recommendations remains underexplored.
+Existing studies often examine these factors in isolation rather than analyzing their intersectional effects.
+Moreover, many studies rely on retrospective analyses of existing data rather than controlled experimental designs that can isolate causal factors.
+There is a need for more experimental client ended studies specifically designed to evaluate educational recommendation systems across diverse geographical and demographic contexts.
 3 Methodology: Evaluation Framework for University Recommendations
@@
 $Acc(s,u)=e^{-\lambda\cdot d(s,u)}$ (1)
-where $\lambda$ is a decay parameter and $d(s,u)$ is the geodesic distance (in km) between the capital cities of the student’s and university’s countries, calculated using Vincenty’s formula Vincenty (1975) via the geopy library, providing approximate structural distance between a student and institution.
-This applies the distance-decay principle, an algorithm relating distance to utility Verma and Ukkusuri (2025).
+where $\lambda$ is a decay parameter and $d(s,u)$ is the geodesic distance (in km) between the capital cities of the student’s and university’s countries, calculated using Vincenty’s formula vincenty_direct_1975 via the geopy library, providing approximate structural distance between a student and institution.
+This applies the distance‑decay principle, an algorithm relating distance to utility verma_what_2025.
 Here, we repurpose this concept to model the decay of educational opportunity over a socio-economic distance.
@@
 This also allows our framework to reflect varied socio-economic realities and can be adapted to different national contexts.
+We set $\lambda$ values such that we could create an exponentially widening accessibility gap between profiles.
+Through sanity checks, we targeted interpretable cutoffs across distance bands: (i) Regional ($d\leq 3{,}000$ km): high-SES profiles should exhibit high accessibility ($Acc>0.7$); (ii) Mid-range ($d\approx 6{,}000$8,0008{,}000 km): mid-SES accessibility should be moderate ($Acc\in[0.2,0.6]$); (iii) Long-haul ($d\geq 12{,}000$ km): low-SES accessibility should be negligible ($Acc<0.2$).
+These $\lambda$ values were then tuned empirically by comparing accessibility disparities between economic proxies at matched distances.
+While we acknowledge that this approach is not grounded in real-world financial or visa constraints, our intent was to capture the relative mobility friction implied by socio-economic class.
 3.1.2 Reputation Alignment
@@
 Any university ranked beyond this threshold, or not ranked at all, receives a reputation score of 0.
-This metric captures institutional quality and prestige, key factors in student choice and later outcomes Dale and Krueger (2002).
+This metric captures institutional quality and prestige, key factors in student choice and later outcomes dale_estimating_2002.
 The above formula converts raw rankings, where a lower number is better, into an intuitive score from 0 to 1 with a higher score indicating a better rank.
-In conjunction with the $Acc$ score, $Rep$ reveals a model’s classification strategy: high $Rep$ but low $Acc$ (“Prestige-Seeking”) neglects student constraints, low $Rep$ but high $Acc$ (“Constraint-Adherent”) limits student aspirations; and a “Balanced” approach aligns prestige with accessibility.
+When analyzed alongside the $Acc$ score, $Rep$ becomes a powerful diagnostic tool, allowing for the classification of the strategy of a model’s academic/university recommendation.
+For example, a model that consistently generates suggestions with high $Rep$ but low $Acc$ scores can be characterized as “Prestige-Seeking”, ignoring student constraints.
+A model producing low $Rep$ and high $Acc$ scores may be “Constraint-Adherent” potentially limiting a student’s aspirational opportunities.
+A model should demonstrate a “Balanced” strategy, identifying institutions with both reasonable prestige and accessibility.
+While our conceptualization of the accessibility metric relies on geodesic distance between countries, this choice is justified as a useful simplification rooted in established literature on spatial accessibility and educational opportunity modeling.
+Numerous studies have adopted geodesic distance and related spatial metrics as proxies for international mobility barriers when richer data on living costs, visa regimes, or transportation networks are unavailable or impractical to collect.
+For instance, spatial2017 show how spatial accessibility measures grounded in distance yield interpretable estimates of opportunity structure in university admission models.
+Similarly, gravity2014 formalize educational migration as flows mediated by geographic ‘friction’, with distance functioning as a primary accessibility constraint.
+edm2020 further justify the use of direct distance metrics in predictive models of student mobility when individual-level factors cannot be exhaustively incorporated.
+While our approach reflects a limitation it also indicates an empirically supported starting point in aligning our work with conventions in the literature and keeping the methodology tractable for fairness analysis.
 3.1.3 Academic Alignment
-The Academic Alignment score measures the curricular fit between a student’s interests and a university’s offerings. It is defined using a formula analogous to a Jaccard index Travieso et al. (2024).
+The Academic Alignment score measures the curricular fit between a student’s interests and a university’s offerings.
+It is defined using a formula analogous to a Jaccard index travieso_analytical_2024.
 $Acad(s,u)=\frac{|T_{s}\cap T_{u}|}{|T_{s}\cup T_{u}|}$ (3)
@@
 A score greater than 1 (clipped to 1.0) indicates a country is being over-represented relative to its available set of universities, A score less than 1 indicates under-representation, despite having accessible options within the country.
-This tackles a key source of bias in global recommender systems: the dominance of countries with large higher education sectors Yi et al.
-(2019).
+This tackles a key source of bias in global recommender systems: the dominance of countries with large higher education sectors yi_sampling-bias-corrected_2019.
 An LLM trained on web data will encounter vastly more text about U.S.
@@
 4.1.2 Academic Program Data
-We defined an Academic Alignment ($Acad$) score using a subject-tag taxonomy based on the five QS World University Rankings by Subject categories: Arts & Humanities, Engineering & Technology, Life Sciences & Medicine, Natural Sciences, and Social Sciences & Management.
+We defined an Academic Alignment ($Acad$) score using a subject‐tag taxonomy based on the five QS World University Rankings by Subject categories: Arts & Humanities, Engineering & Technology, Life Sciences & Medicine, Natural Sciences, and Social Sciences & Management.
 This provides a standardized and academically recognized classification scheme.
-Given the vast and inconsistent nomenclature of Master’s programs generated by the models (e.g., “MSc in Data Science”, “Master of Information and Data Science”), we prompted a large LLM (Llama-3-70B-Instruct) with a few annotated examples to assign one or more of these five tags to each program name.
+Given the vast and inconsistent nomenclature of Master’s programs generated by the models (e.g., “MSc in Data Science”, “Master of Information and Data Science”), we prompted a large LLM (Llama‑3‑70B‑Instruct) with a few annotated examples to assign one or more of these five tags to each program name.
 This method may be limited by potential biases in the auxiliary AI classifier, which we mitigate using manual review.
@@
 4.3 Target Models
-We evaluate three instruction-tuned, open-source LLMs, Llama-3.1 AI (2024), Gemma Team et al.
-(2024), and Mistral Jiang et al.
-(2023), chosen for their research popularity, open accessibility (vital for reproducibility), and diverse origins (Meta, Google, Mistral AI).
+We evaluate three instruction‑tuned, open‑source LLMs, Llama‑3.1 meta_llama31_2024, Gemma gemma_team_2024, and Mistral jiang2023mistral, chosen for their research popularity, open accessibility (vital for reproducibility), and diverse origins (Meta, Google, Mistral AI).
 Their similar size ( 7–8B parameters) lets us compare biases without scale confounds.
-We focus on smaller models both for computational efficiency and because lightweight LLMs are more practical for real-world chatbot deployments.
+We focus on smaller models both for computational efficiency and because lightweight LLMs are more practical for real‑world chatbot deployments.
 4.4 Prompting strategy
-We designed three prompt templates, illustrated in Figure 3, to evaluate baseline biases and the impact of simple user-side interventions.
+We designed three prompt templates, illustrated in Figure 3, to evaluate baseline biases and the impact of simple user‑side interventions.
 Table 1: Demographic Representation Score and its components for the base prompt, by demographic factors
@@
 The base template is a standard university recommendation query with demographic placeholders.
-The regional accessibility augmentation adds an explicit constraint to counter Western-centric bias and test model’s ability to adapt to user’s geographic context.
+The regional accessibility augmentation adds an explicit constraint to counter Western‑centric bias and test model’s ability to adapt to user’s geographic context.
 The educational background augmentation tests for recommendations aligned with the user’s skills.
 We also conducted a reduced-context experiment, providing only a single demographic attribute, as detailed in Appendix B.
-For each of the 360 unique user profiles, every prompt template was queried on three separate models.
-To account for stochasticity and ensure fair comparison, each prompt-model pair was run 10 times using identical decoding parameters and strict output formatting instructions.
+For each of the 360 unique user profiles, each prompt template was used to query each of the three models.
+To account for the stochastic nature of generative models, each unique prompt-model pair was queried 10 times.
+This resulted in a total of 21,600 responses.
+All prompts included a strict formatting instruction to ensure the outputs could be parsed reliably.
+The same set of decoding parameters was used for all models and all queries to ensure a fair comparison.
 4.5 Implementation
@@
 Crucially, major developing nations like India and Brazil still scored GRS=0 across all models, even under regional constraints.
-Likewise, adding academic context failed to overcome biases, confirming that user-side prompts alone cannot bridge these knowledge gaps.
-Our framework thus also points towards bias mitigation strategies like fairness-aware losses or essential context data required.
-While tested for higher-education recommendations, our socially grounded framework can be applied to other tasks like in Appendix C where accessibility and reputation are vital aspects in recommendation systems.
+Likewise, adding academic context failed to overcome biases, confirming that user‑side prompts alone cannot bridge these knowledge gaps.
+Our framework thus also points towards bias mitigation strategies like fairness‑aware losses or essential context data required.
+While tested for higher‑education recommendations, our socially grounded framework can be applied to other tasks like in Appendix C where accessibility and reputation are vital aspects in recommendation systems.
 6 Conclusion
-This paper delivers a comprehensive analysis of how open-source LLMs shape higher-education recommendations, uncovering stark demographic and geographic biases.
-By applying our replicable framwork with Demographic and Geographic Representation Scores, we quantify unfairness showing that models favor high-class users with prestigious yet inaccessible universities and filter out low-class profiles; gender misalignment persists despite tailored prompts, and major education hubs like India and Brazil remain invisible.
-Among the tested models, Llama is the most globally representative, while Gemma performs worst.
-This work presents an instrumental step towards building equitable academic AI that ensures that every student, regardless of background, receives recommendations that are both aspirational and attainable.
+This paper delivers a comprehensive analysis of how open‑source LLMs shape higher‑education recommendations and exposes the stark biases they encode.
+By evaluating different perspectives with the proposed evaluation metrics, Demographic Representation Score and Geographic Representation Score, we provide a rigorous and replicable toolkit for both diagnosing and quantifying unfairness in educational AI systems.
+Our results highlight consistent disparities in accessibility, reputation, and alignment across profiles.
+Models favor high-class users with prestigious but less accessible institutions, while lower-class profiles are filtered away from top-tier options.
+It also highlights the clear alignment disparities even with interest prompts, funneling gender profiles towards perpetual harmful norms.
+Through all this, the fundamental problem that presents itself is the lack of global representation with major higher education hubs still overshadowed, signaling a profound blind spot.
+Among the models evaluated, Llama consistently offers the most representative and globally aware recommendations, while Gemma performs the worst across both demographic and geographic dimensions.
+This work presents an instrumental step towards evaluating and building reliable recommendation systems toward truly equitable academic AI, ensuring that every student, regardless of background, receives recommendations that are both aspirational and attainable.
 7 Limitations
-While this analysis represents a comprehensive examination of bias in LLM-based academic recommendation, there are a few limitations to be considered:
+While this analysis represents a comprehensive examination of bias in LLM‑based academic recommendation, there are a few limitations to be considered:
   • Synthetic Profile Scope.
-Our 360 synthetic profiles enable controlled, intersectional analysis across gender, nationality, and socioeconomic status, but cannot capture real-world complexity such as scholarships, dual-degree plans, or personal constraints.
+Our 360 synthetic profiles enable controlled, intersectional analysis across gender, nationality, and socioeconomic status, but cannot capture real‐world complexity such as scholarships, dual‑degree plans, or personal constraints.
   • Dependence on QS Rankings.
 Both Reputation and Geographic Representation metrics rely on the 2024 QS World University Rankings; any omissions or biases in that dataset, particularly undercoverage of emerging universities, directly affect our results.
-  • Subject-Tag Taxonomy Reliance.
+  • Subject‑Tag Taxonomy Reliance.
 Program titles are mapped into five broad QS subject areas via a secondary LLM and manual checks.
@@
   • Model and Scale Constraints.
-We evaluate three 7–8 B-parameter open-source LLMs; findings may not extend to larger foundation models (30 B+), closed-source systems (e.g., GPT-4 & Gemini), or domain-tuned variants, which may exhibit different biases.
+We evaluate three 7–8 B‑parameter open‑source LLMs; findings may not extend to larger foundation models (30 B+), closed‑source systems (e.g., GPT‑4 & Gemini), or domain‑tuned variants, which may exhibit different biases.
   • Fixed Decay Parameters.
@@
 Our evaluation framework goes a step further than standard metrics by providing different perspectives for practitioners to understand what a model lacks.
-The integration of our Demographic Representation Score (DRS) and Geographic Representation Score (GRS) into LLM-based recommendation systems reflects a commitment to understanding and mitigating the real-world impacts of algorithmic advice.
+The integration of our Demographic Representation Score (DRS) and Geographic Representation Score (GRS) into LLM‑based recommendation systems reflects a commitment to understanding and mitigating the real‑world impacts of algorithmic advice.
 Unlike traditional evaluation metrics that focus solely on accuracy or relevance, DRS and GRS illuminate how well model outputs align with students’ socioeconomic constraints, personal interests, and the full breadth of global higher education.
 Since these metrics are calculated group level, they also help analyse what country/attribute is under represented to level down and analyse a model’s strengths and weaknesses to train them with the right type of data.
-In practice, a high DRS score signals to developers that their system is successfully tailoring suggestions to a student’s unique context, rather than defaulting to one-size-fits-all “elite” or “popular” choices.
-Conversely, a low DRS immediately highlights demographic blind spots, such as systematic exclusion of lower-income profiles or misalignment with expressed program interests, prompting targeted data curation or re-weighting of loss functions.
+In practice, a high DRS score signals to developers that their system is successfully tailoring suggestions to a student’s unique context, rather than defaulting to one‑size‑fits‑all “elite” or “popular” choices.
+Conversely, a low DRS immediately highlights demographic blind spots, such as systematic exclusion of lower‑income profiles or misalignment with expressed program interests, prompting targeted data curation or re‑weighting of loss functions.
 Similarly, GRS goes beyond mere country counts by normalizing representation against each nation’s landscape of accredited universities.
-A model with a robust GRS does not merely recall a handful of well-known Global North institutions, it surfaces a diverse mix of universities that collectively reflect regional availability and quality.
-Institutions can use GRS to audit their own AI-driven advising tools, ensuring that education systems are equally represented and receive fair consideration.
+A model with a robust GRS does not merely recall a handful of well‑known Global North institutions, it surfaces a diverse mix of universities that collectively reflect regional availability and quality.
+Institutions can use GRS to audit their own AI‑driven advising tools, ensuring that education systems are equally represented and receive fair consideration.
 Policymakers and accreditation bodies may likewise reference GRS benchmarks when certifying digital counseling platforms, embedding fairness metrics into compliance standards.
@@
 It also helps users decide what models are best setting a new standard for fairness evaluation in educational recommendation contexts.
-Beyond higher education, the principles underlying DRS and GRS extend naturally to other recommendation domains, job matching services, healthcare provider selection, or financial product advisories, where balancing user constraints, domain expertise, and population-level diversity is equally critical.
+Beyond higher education, the principles underlying DRS and GRS extend naturally to other recommendation domains, job matching services, healthcare provider selection, or financial product advisories, where balancing user constraints, domain expertise, and population‐level diversity is equally critical.
 Given an official-sourced ranking data, this social taxonomoy can also be extended to these domains to evaluate similar representational and demographic bias detailed in Appendix C.
-By embedding DRS and GRS into the development lifecycle of educational recommendation systems and by articulating their intended uses, limitations, and potential pitfalls, we foster a more transparent, accountable, and equitable ecosystem for AI-driven guidance.
-Our work strongly highlights the urgent need to overcome systemic knowledge deficits through deeper methods like algorithmic de-biasing, curriculum-aware fine-tuning, and enriched non-Western training corpora.
+By embedding DRS and GRS into the development lifecycle of educational recommendation systems and by articulating their intended uses, limitations, and potential pitfalls, we foster a more transparent, accountable, and equitable ecosystem for AI‑driven guidance.
+Our work strongly highlights the urgent need to overcome systemic knowledge deficits through deeper methods like algorithmic de‑biasing, curriculum‑aware fine‑tuning, and enriched non‑Western training corpora.
 Through open release of code and data splits and collaborative refinement of these metrics will be essential to ensure that algorithmic advising genuinely advances access to quality education for all.
@@
 We are grateful to Mr.
-Krishnan Narayanan (Ihihaasa/IIT Madras), Prof.
+Krishnan Narayanan (itihaasa/IIT Madras), Prof.
 Vanessa Evers (University of Twente, Netherlands), Dr.